# Inductive Graph Representation Learning with Quantum Graph Neural Networks

### v1

@@ 4.1 Inner Workings of GraphSAGE
In GNNs, embedding generation consists of two main steps: sampling and aggregation performed by GCNs. In its layer $l=0$, the message of target node $v$ is $\mathbf{h}_{v}^{(0)}=\mathbf{x}_{v}$, where $\mathbf{x}_{v}$ represents the input features of $v$. Using the computational graph representation of the input graph, the sampling step collects features from neighboring nodes $u$ of $v$. The message (Msg) generated at the GCN layer $l$ for the neighbor $u$ of $v$ reads
@@
where $\mathcal{N}(v)$ represents the set of neighbors of node $v$. The embeddings $\mathbf{h}_{u}^{(l-1)}$ correspond to the representations of neighbor $u$ from the previous layer, $l-1$. In the message-passing process, learnable parameters $\mathbf{W}^{(l)}$ (weights) and $\mathbf{B}^{(l)}$ (biases) are assigned at layer $l$ to the neighboring nodes $u$ and the target node $v$, respectively. In the aggregation step, neighbor messages are combined as follows
@@
In contrast, GraphSAGE introduces an intermediate step, where $\mathbf{h}_{v}^{(l)}$ is not updated automatically. In fact, it first enables a flexible aggregation process to generate the aggregated embedding of neighboring nodes $\mathbf{h}_{\mathcal{N}(v)}^{(l)}$, typically expressed as
where Agg can be any general vector function. Common aggregation methods include Mean, LSTM [44], or Pooling [45]. In this work, we choose the Mean function due to its simplicity and permutation invariance. This type of aggregator computes the element-wise mean of $\{\mathbf{h}_{u}^{(l-1)}|\,u\in\mathcal{N}(v)\}$, providing a simple yet effective approach to combining neighbor information. Hence, we have $\mathbf{h}_{\mathcal{N}(v)}^{(l)}=\textsc{mean}(\{\mathbf{h}_{u}^{(l-1)},\forall u\in{\mathcal{N}(v)}\}$, which closely resembles the aggregation procedure in the transductive GCN framework [15]. The second key feature of GraphSAGE is the concatenation of the target node’s previous-layer embedding $\mathbf{h}^{(l-1)}_{v}$ with $\mathbf{h}^{(l)}_{\mathcal{N}(v)}$, which is not immediately done in GCNs. Now, the final embedding of $v$ reads
$\mathbf{h}_{v}^{(l)}=\sigma\left(\mathbf{W}^{(l)}\cdot\textsc{Concat}\left(\mathbf{h}_{\mathcal{N}(v)}^{(l)},\mathbf{h}_{v}^{(l-1)}\right)\right).$ (4)
where $\sigma$ is a nonlinear activation function that enhances expressivity. The function Concat denotes the concatenation of embeddings. Alternatively, some GNN architectures refine this process by integrating attention mechanisms, which dynamically adjust the influence of each neighbor based on its relevance [46].
In the layer $L$, the final node embeddings, $\mathbf{z}_{v}=\mathbf{h}_{v}^{(L)}$ encodes both the target node’s intrinsic features and its structural relationships within the graph. Hence, relying fundamentally on the features, GraphSAGE effectively captures the structural information of the graph by learning the role of a node within it. Building upon these foundations, we now extend this framework to the quantum domain, introducing the Quantum Graph Neural Network.
5 Proposed QGNN Framework
In this section, we present our QGNN framework, which integrates quantum circuits with GraphSAGE techniques to generate node embeddings. The QGNN framework comprises two key components: (1) Use of aggregation and concatenation mechanisms to generate embeddings from a node’s local neighborhood features. (2) A quantum circuit that acts as a modular aggregator, combining the embeddings of connected nodes.
The first component, inspired by GraphSAGE, utilizes the adjacency matrix of the input graph to construct its computation graph representation, preserving local graph structures. This enables the identification of connected nodes, whose features are sampled uniformly [^1] Following Ref. [23], we uniformly sample neighbors $u$ from $\{u\in\mathcal{V}:(u,v)\in\mathcal{E}\}$. and inputted into an MLP (see Fig. 1(b)). The second component utilizes Quantum Graph Convolutional Networks (QGCNs) as a replacement for classical MLPs to perform GraphSAGE’s aggregation. This approach leverages quantum mechanics to process node information, as illustrated in Fig. 2(a). The input data is passed to the quantum circuit through a feature map, which encodes classical data into an $n$-qubit input state $\rho_{\text{in}}$, spanning an $n$-qubit Hilbert space $\mathcal{H}$. Concatenation remains unchanged, as it is a purely vector-based operation within this framework that occurs in the classical post-processing phase.
A key implication of the QGCN design is that, for nodes with feature vectors of identical dimension, the number of qubits required for encoding depends solely on this dimension. By sequentially inputting node feature vectors into the circuit, the architecture remains invariant to the number of sampled nodes and edges. Importantly, the number of qubits can be increased without altering the number of features per node, enabling additional qubits per feature for encoding. However, the minimum number of qubits must still match the number of features per node.
Together, these elements constitute our QGNN framework for node-level tasks. Fig. 2 outlines the tools and building blocks required to construct the complete QGNN. This approach extends inductive learning for graph-structured data into the quantum domain, offering a versatile and scalable framework.
(a) QGNN framework
(b) Building blocks for QGCNs.
Figure 2: Schematic representation of the proposed QGNN framework and the fundamental building blocks for constructing the QGCN circuit. (a) Embedding generation process: Based on the computational graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, node features of connected nodes (red and yellow) are sampled and input into the QGCN circuit. The QGCN aggregates these features to create neighboring node embeddings (orange), which are concatenated with the target node’s self-embedding (purple) to produce its updated embedding. To enhance expressivity, the concatenated representation undergoes a non-linear transformation $\sigma$. (b) Abstract description of operations for constructing the quantum convolutional (C) and quantum pooling (P) layers of the QGCN circuit. A feature map (FM) encodes classical inputs into an $n$-qubit state.
5.1 The QGCN circuitry
Inspired by previous QCNN designs [24], the QGCN also alternates between quantum convolutional and quantum pooling layers.
Convolutional layers encode local graph structures by applying two qubit unitaries $W_{ij}$ on pairs of neighboring qubits $i$ and $j$. These unitaries utilize parameterized quantum gates, such as $R_{X}$ and $R_{Z}$ with tunable angles, which constitute the Ansatz. Controlled gates such as CZ and CNOT entangle parametrized states across qubits. This allows the quantum circuit to refine embeddings, improving expressivity while effectively capturing local structures.
Pooling layers reduces the dimensionality of quantum states while retaining essential features. This reduction simplifies data representation, decreases computational complexity, while retaining key information. Pooling is achieved by measuring a subset of qubits and using the measurement outcomes to apply classical-conditioned Pauli gates on adjacent qubits. However, the principle of deferred measurement [47] is applied as a design choice, and hence the pooling layers employ controlled operations, with (partial) measurements postponed at the end of the computation.
A generic QGCN circuit is described as a composition of quantum channels as
$\Phi_{\bm{\theta},\bm{\lambda}}=\bigcirc_{l=1}^{L}\big{(}\text{P}_{l}^{\bm{\lambda}_{l}}\circ\text{C}_{l}^{\bm{\theta}_{l}}\big{)},$ (5)
where $\circ$ denotes a single function composition, and $\bigcirc_{l=1}^{L}$ represents the composition of $L$ functions applied sequentially. Each QGCN layer $l$ comprises a quatnum convolutional layer $\text{C}_{l}^{\bm{\theta}_{l}}$ followed by a quantum pooling layer $\text{P}_{l}^{\bm{\lambda}_{l}}$, with $\bm{\theta}_{l}$ and $\bm{\lambda}_{l}$ being the convolution and pooling parameters, respectively. The alternating structure of the QGNN circuit processes and simplifies the input quantum state, starting with $\text{C}_{1}$ and ending with $\text{P}_{L}$.
The convolutional layer $\text{C}_{l}^{\bm{\theta}_{l}}:\mathcal{S}(\mathcal{H}_{l})\rightarrow\mathcal{S}(\mathcal{H}_{l})$ preserves the size of the quantum register. It reads
@@ 5.1 The QGCN circuitry
Figure 3: Schematic representation of the QGCN circuit. The QGCN takes as input an $n$-qubit quantum state $\rho_{\text{in}}$ and processes it through a sequence of $L$ alternating convolutional (C) and pooling (P) layers. In each convolutional layer, parameterized quantum gates perform unitary transformations $W$ on pairs of neighboring qubits. The pooling layers then reduce the number of qubits by tracing out half of them between layers, preserving essential features while compressing the state space dimensionality. After the last pooling layer, the circuit produces an output state $\rho_{\text{out}}$, over which the expectation value of an operator $\mathcal{O}$ is measured.
The pooling layer $\text{P}_{l}^{\bm{\lambda}_{l}}:\mathcal{S}(\mathcal{H}_{l})\rightarrow\mathcal{S}(\mathcal{H}_{l+1})$ reduces the size of the quantum register by tracing out specific qubits, such that $\dim(\mathcal{H}_{l+1})<\dim(\mathcal{H}_{l})$, and is defined as
[…]
where the $i$-qubit is traced out in the $l$-th layer. See Fig. 2(b)) for a schematic representation. The pooling layers typically discard half of the qubits at each step. The complete QGCN circuit is illustrated in Fig. 3.
We adopt a simple architecture for $\text{C}_{l}^{\bm{\theta}_{l}}$ and $\text{P}_{l}^{\bm{\lambda}_{l}}$, as minimal designs often achieve comparable or superior results compared to more complex counterparts [33]. In this context, the unitary $W$ is defined as in Fig. 4, which is an optimal decomposition in the number of gates according to Ref. [48]. This minimal design employs the smallest number of quantum operations necessary to prepare any two-qubit unitary (up to a global phase), requiring only three CNOTs, three single-qubit rotation, and four general single-qubit $R^{G}$ gates. We define the general rotation $R^{G}$ in layer $l$ as $R_{G}(\bm{\theta}_{l})=e^{-iX\theta^{1}_{l}/2}e^{-iZ\theta^{2}_{l}/2}e^{-iX\theta^{3}_{l}/2}$. Generally, entanglement is achieved applying non-parametrized CZ gates [^4] In this context, CNOTs are replaced with CZs in Fig. 3.. In the pooling layers, entanglement is followed by local measurements on $\mathcal{O}$, enabled by the deferred measurement principle [47]. The chosen design reduces the complexity of the QGCN circuit (5) by
$\Phi_{\bm{\theta},\bm{\lambda}}\to\Phi_{\bm{\theta}}.$ (8)
[…]
The convolutional cell $W$ is customizable, and its definition uniquely determines the convolutional layers. In particular, QGCN layers can preserve data symmetries [49, 50, 51], making them, for example, particularly effective for symmetry-equivariant tasks. The circuit can also be extended to include QEC protocols to incorporate fault-tolerant computation [24, 52].
We now proceed to benchmark our proposed framework on the QM9 dataset, comparing its performance against its classical counterpart in predicting molecular properties.
6 Results
6.1 Implementation details
We now present the training and testing results of our QGNN framework on the QM9 dataset, comparing the performance of classical GNNs with our quantum design. Performance is evaluated using the Smooth L1 Loss and the R2 score, with an initial focus on a small subset of molecular samples. The selected Loss is chosen for its stability and robustness to outliers, making it particularly suitable for small sample sizes with high variability. Additionally, we employ the Adam Optimizer [53] with $\beta_{1}=\beta_{2}=0.9$ and a decreasing learning rate.
We train our QGNN in a supervised learning manner to predict the energy of the lowest unoccupied molecular orbital ($E_{\text{LUMO}}$). The algorithm processes molecules in two steps: (1) atom-level aggregation, which combines atom (or node) features with neighborhood information, and (2) molecule-level aggregation, which pools atom information into a single molecular representation. The loss function compares the predicted results with the target energy. In that way, we ensure that local and global structural information is effectively captured.
In the proposed algorithm, we consider the following: first, we use all possible hops $k$ for each molecule in both classical and quantum benchmarking, where $k_{\text{max}}$ equals the number of atoms in the molecule being processed. For instance, a 9-atom molecule undergoes 9 hops, propagating embeddings from lower-index to higher-index atoms. This restricts the computational graph to a single branch starting from the smallest-index atom. Second, this design choice eliminates the need for a final QNN layer to process multiple target nodes. This minimal approach significantly reduces computational overhead, enabling us to focus solely on benchmarking the (Q)GCNs. In addition, to avoid overfitting when aggregating the atom embeddings in both scenarios, we employ a single (Q)GCN model across the (Q)GNN framework. This ensures consistency and reduces the risk of overfitting given the small sample sizes. For more details on the algorithm, refer to App. A.
The classical GNN implemented in PyTorch [^5] For classical GNN implementations, we construct a custom GCN from scratch using PyTorch, ensuring closer alignment with our QGNN implementation and enabling a fair classical-quantum comparison[54] consists of a single hidden layers of 8 neurons, respectively. The QGCN circuit employes an 8-qubit  [^6] The number of qubits equals the node feature dimension plus one (accounting for concatenation with the mean of the target node’s embedding). Refer to App. A for more information., implemented with Qadence [55]. In this setup, the QGCN has substantially fewer tunable parameters (113) compared to its classical counterpart (209) - less than half the total parameters. Classical data from the QM9 dataset is encoded into the quantum circuit using a Fourier Feature Map [41, 56]. The QGNN comprises of three quantum convolutional layers with repetitions of $r_{1}=r_{2}=r_{3}=1$, and two quantum pooling layers, as defined in Eq. (7). Finally, measurements evaluate the local magnetization, defined as the expectation value of the $Z$ Pauli operator, $\hat{M}=\sum_{i}\sigma_{z}^{(i)}/N$.
This aforementioned setup and metaparameters, along with the chosen loss and R² score metrics, will be employed consistently throughout the results section unless explicitly stated otherwise.
6.2 Comparison Between Classical and Quantum Models

### v2

@@ IV.1 Inner Workings of GraphSAGE
In the layer $l=0$, the embedding of the target node $v$ is $\mathbf{h}_{v}^{(0)}=\mathbf{x}_{v}$, where $\mathbf{x}_{v}$ represents the input features of $v$. Based on the computational graph, the sampling step uniformly selects the neighbors $u$ of node $v$ and retrieves their feature vectors. The message (Msg) generated at the GCN layer $l$ for the neighbor $u$ of $v$ reads
@@
where $\mathcal{N}(v)$ represents the set of neighbors of node $v$. The embeddings $\mathbf{h}_{u}^{(l-1)}$ correspond to the representations of neighbor $u$ from the previous layer, l1l-1. In the message-passing process, learnable parameters $\mathbf{W}^{(l)}$ (weights) and $\mathbf{B}^{(l)}$ (biases) are assigned at layer $l$ to the neighboring nodes $u$ and the target node $v$, respectively. In the aggregation step, neighbor messages are combined as follows
@@
being the aggregation $(\textsc{Agg})$ originally defined as the average of the neighboring nodes’ embeddings. In contrast, GraphSAGE introduces an intermediate step, where $\mathbf{h}_{v}^{(l)}$ is not updated automatically. In fact, it first enables a flexible aggregation process to generate the aggregated embedding of neighboring nodes $\mathbf{h}_{\mathcal{N}(v)}^{(l)}$, typically expressed as
where Agg can be any general vector function. Common aggregation methods include Mean [welling_GCN_2017], LSTM [hochreiter_1997_long], or Pooling [qi_2016_pointnet]. In this work, we choose the Mean function due to its simplicity and permutation invariance, i.e., $\mathbf{h}_{\mathcal{N}(v)}^{(l)}=\textsc{mean}(\{\mathbf{h}_{u}^{(l-1)},\forall u\in{\mathcal{N}(v)}\}$. This type of aggregator computes the element-wise mean of the neighbor embeddings $\mathbf{h}_{u}^{(l-1)}$. GraphSAGE also performs concatenation of the target node’s previous-layer embedding $\mathbf{h}^{(l-1)}_{v}$ with $\mathbf{h}^{(l)}_{\mathcal{N}(v)}$, which is not immediately done in [welling_GCN_2017]. The final embedding of $v$ reads
$\mathbf{h}_{v}^{(l)}=\sigma\left(\mathbf{W}^{(l)}\cdot\left[\mathbf{h}_{\mathcal{N}(v)}^{(l)}\,\|\,\mathbf{h}_{v}^{(l-1)}\right]\right)$ (4)
where $\cdot\|\cdot$] denotes concatenation and $\sigma$ is a nonlinear activation function that enhances expressivity. Alternatively, some GNN architectures refine this process by integrating attention mechanisms, which dynamically adjust the influence of each neighbor based on its relevance [petar_gats_2018].
In the layer $L$, the final node embedding, $\mathbf{h}_{v}^{(L)}$ encodes both the target node’s intrinsic features and its structural relationships within the graph. Hence, GraphSAGE effectively captures the structural information of the graph by learning the role of a node within it. Building upon these foundations, we now extend this framework to the quantum domain, introducing our QGNNs.
V Proposed QGNN Framework
In this section, we present our QGNN framework, which integrates quantum circuits with GraphSAGE techniques to generate node embeddings. The QGNN framework comprises three key components: (1) the updated target node embedding is generated by aggregating and concatenating messages from neighboring nodes. (2) A feature map that maps classical features into qubit states. (3) A QGCN circuit that acts as the quantum modular aggregator, combining the embeddings of connected nodes.
The first component, inspired by GraphSAGE, utilizes the adjacency matrix of the input graph to construct its computation graph representation, preserving local graph structures. This enables the identification of connected nodes $u$ of $v$, i.e., $\{u\in\mathcal{V}:(u,v)\in\mathcal{E}\}$, whose features are sampled uniformly and input into an MLP (see Fig. 1(b)). The second component initializes the qubit states, while the third employs QGCNs to perform GraphSAGE-style aggregation, replacing the classical MLPs traditionally used in the message-passing phase. This approach leverages quantum mechanics to process local classical node features. Together, these elements form our QGNN architecture for node-level tasks. Fig. 3 illustrates the overall framework.
Figure 2: QGNN framework.
Figure 3: Schematic representation of the proposed QGNN framework. Embedding generation process: based on the computational graph of $\mathcal{G}=(\mathcal{V},\mathcal{E})$, node features of connected nodes to $v$ (red and yellow) are sampled uniformly and input into the QGCN. The circuit aggregates these features to create a single, modular neighboring-node embedding (orange) $\mathbf{h}_{\mathcal{N}(v)}^{(L)}(\bm{\theta}_{v})$. This is concatenated classically with the target node’s self-embedding (purple) through a simple vector operation and then processed by a non-linear function $\sigma$ to yield the updated embedding $\mathbf{h}_{v}^{(L)}(\bm{\theta}_{v})$.
V.1 Quantum feature map
The classical input feature of node $u$, denoted by $\mathbf{x}_{u}$, is passed into the QGCN through a feature map. The feature map $\mathcal{F}$ is realized by a unitary $U_{x}$ acting on the initial $n$-qubit state $|0\rangle^{\otimes N}$, spanning the Hilbert space $\mathcal{H}$. Concretely,
$\mathbf{x}_{u}$ $\mapsto|\mathcal{F}(\mathbf{x}_{u})\rangle=U_{x}|0\rangle^{\otimes N},$ (5)
Hence, the initial state of $u$, $\rho_{u}$, becomes
$\rho_{u}=\lvert\mathcal{F}(\mathbf{x}_{u})\rangle\langle\mathcal{F}(\mathbf{x}_{u})\rvert$. (6)
In this particular work, we employ the Fourier quantum feature map to improve expressivity, as it effectively maps highly nonlinear classical functions into qubit states [kyriienko_2021_solving, schuld_2021]. For a set of $N$ neighbors of node $v$, the overall input state becomes
$\rho_{\mathrm{in}}=\bigotimes_{w=1}^{N}\,\rho_{w},$ (7)
Eq. (7) implies that each state $\rho_{w}$ has the same dimension, since the number of qubits defining it depends solely on the dimension of $\mathbf{x}_{u}$ (see Eq. (6)), which is assumed to be identical for all vectors. By feeding node features into the circuit sequentially, the architecture remains independent of the number of sampled nodes and edges. The number of qubits may be increased to allow richer encodings, but the minimum must match the feature dimension. Now, let us describe the Ansatz.
V.2 The QGCN circuitry
Inspired by previous QCNN designs [lukin_qcnn_2019], the QGCN circuit also alternates between quantum convolutional and quantum pooling layers and defines the local Ansatz.
Convolutional layers capture local graph structure by applying learnable two-qubit operations that generate entanglement between neighbouring qubits. This produces quantum states that encode the relevant local features. Conversely, pooling layers reduce the dimensionality of the quantum state while retaining its essential features, thereby lowering computational cost without losing key information. In this way, pooling yields a coarser quantum representation that complements the local state transformations performed by the convolutional layer.
The QGCN circuit consists of a composition of $L$ quantum channel layers, defined as
$\Phi_{\bm{\theta}_{v}}=\bigcirc_{l=1}^{L}\left(\mathrm{P}_{l}\circ\mathrm{C}_{l}^{\bm{\theta}^{l}_{v}}\right),$ (8)
where $\bm{\theta}_{v}:=\{\bm{\theta}^{l}_{v}\}_{l=1}^{L}$. We denote a single function composition by $\circ$, while $\bigcirc_{l=1}^{L}$ refers to the sequential composition of $L$ functions. Importantly, each QGCN layer $l$ is built from a quantum convolution layer $\mathrm{C}_{l}^{\bm{\theta}^{l}_{v}}$, parametrized by $\bm{\theta}_{v}^{l}$, followed by a quantum pooling layer $\mathrm{P}_{l}$. Since only the convolutional layers carry parameters, they define the Ansatz. Pooling layers could also be parametrized, e.g. $\mathrm{P}_{l}^{\bm{\lambda}_{l}}$, to allow more flexible pooling operations, but we keep the architecture minimal to emulate the quantum analogue of the classical mean pooling. The alternating sequence of convolution and pooling layers progressively transforms and reduces the quantum state, beginning with $\mathrm{C}_{1}$ and ending with $\mathrm{P}_{L}$.
The convolutional layer $\mathrm{C}_{l}^{\bm{\theta}_{v}^{l}}$: $\mathcal{S}(\mathcal{H}_{l})\rightarrow\mathcal{S}(\mathcal{H}_{l})$ preserves the size of the quantum register. It reads
@@ V.2 The QGCN circuitry
Within each convolutional layer, which acts on an even number of qubits, the unitary $W_{l}$—referred to as the convolutional cell—operates on pairs of neighboring qubits $(i,i+1)$, as depicted in Fig. 4. For $n_{l}$ qubits in layer $l$, the operator $W_{l}$ acts on adjacent qubit pairs in an alternating pattern: for even values of $j$, it targets the pairs $(0,1),(2,3),\dots$, while for odd $j$, it instead acts on $(1,2),(3,4),\dots$. Formally, the set of qubit pairs involved are given by $S(j)=\{(i,i+1)\,\mid\,i\equiv j\pmod{2},\,0\leq i\leq n_{l}-2\}$. The parameter $r$ specifies the depth of each convolutional layer $\mathrm{C}_{l}^{\bm{\theta}_{v}^{l}}(\cdot)$. For example, in a two-layer QGCN architecture, the depths of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are denoted as $r_{1}$ and $r_{2}$, respectively. This is $r=[r_{1},r_{2}]$.
Following Ref. [vatan_2004_optimal], $W_{l}$ is generally decomposed into 3 CZ gates, 3 single-qubit rotations, and 4 general $R^{G}$ gates, where
$R^{G}(\bm{\theta}_{v}^{l})=e^{-iX\theta_{1}^{l}/2}e^{-iZ\theta_{2}^{l}/2}e^{-iX\theta_{3}^{l}/2}.$ (10)
The unitary $W$ is defined as in Fig. 4. The convolutional cell $W$ is customizable, and its definition uniquely determines the convolutional layers. In particular, QGCN layers can be designed to preserve data symmetries [larocca_2022_group, chinzei_2023, cerezo_2022_theory_equivariant], making them particularly effective for symmetry-equivariant tasks. The circuit can also be extended to include fault-tolerant operations on QEC codes [lukin_qcnn_2019, zeng_2016].
Figure 5: Schematic representation of the QGCN circuit. The QGCN takes as input an $n$-qubit quantum state $\rho_{\text{in}}$ and processes it through a sequence of $L$ alternating convolutional (C) and pooling (P) layers. In each convolutional layer, parameterized quantum gates perform unitary transformations $W$ on pairs of neighboring qubits. The pooling layers then reduce the number of qubits by tracing out half of them between layers. After the last pooling layer, the circuit produces the output state $\rho_{\mathrm{out}}^{\bm{\theta}_{v}}$.
Conversely, the pooling layer $\mathrm{P}_{l}$: $\mathcal{S}(\mathcal{H}_{l})\rightarrow\mathcal{S}(\mathcal{H}_{l+1})$ reduces the size of the quantum register by tracing out half of the qubits, such that $\dim(\mathcal{H}_{l+1})<\dim(\mathcal{H}_{l})$. It reads
[…]
Here, $i$ denotes the qubits traced out at layer $l$, which halves the quantum register at each pooling step, i.e., $i=\{j\in\{0,1,\dots,n_{l}-1\}\,|\,j\equiv 0\pmod{2}\}$. Typically, the the layer is implemented by measuring a subset of qubits, whose classical outcomes condition the application of controlled Pauli gates on their neighbors. Our design follows, however, the principle of deferred measurement [nielsen]. Hence, the pooling layers first employ quantum controlled operations, deferring the (partial) measurements to the end.
At the final QGCN layer $L$, the modular output state produced by the circuit, $\rho_{\mathrm{out}}^{\bm{\theta}_{v}}$, is defined as
$\rho_{\mathrm{out}}^{\bm{\theta}_{v}}$ $=\Phi_{\bm{\theta}_{v}}(\rho_{\mathrm{in}})=\bigotimes_{w=1}^{N}\,\Phi_{\bm{\theta}_{v}}(\rho_{w})$ (12)
The complete QGCN circuit is illustrated in Fig. 5. Note that the CZ gates employed to entangle qubit states, both in the pooling layers and in the convolutional cell $W_{l}$ definition, can be easily replaced by other two-qubit operations, such as CNOTs.
V.3 Embedding generation and model training
After employing the trainable FM to initialize the qubit states and using the QGCN as the quantum aggregator, the classical embedding for the neighbors $u$ of $v$ is obtained through measurements of an observable $\mathcal{O}$}, expressed as
$\mathbf{h}_{\mathcal{N}(v)}^{(L)}(\bm{\theta}_{v})$ =Tr[$=\Tr[\rho_{\mathrm{out}}^{\bm{\theta}_{v}}{\mathcal{O}}]=\langle{\mathcal{O}}\rangle_{\rho_{\mathrm{out}}^{\bm{\theta}_{v}}}.$ (13)
Note that the embedding is generated by the same QGCN regardless of the number of neighboring nodes $u$. It can be further concatenated with the self-embedding of node $v$ to yield its updated representation as
$\mathbf{h}_{v}^{(L)}(\bm{\theta}_{v})=\sigma\left(\left[\mathbf{h}_{\mathcal{N}(v)}^{(L)}(\bm{\theta}_{v})\,\|\,\mathbf{h}_{v}^{(1)}\right]\right).$ (14)
As the QGCN circuit has $L$ layers, the initial embedding $\mathbf{h}_{v}^{(1)}$ can be only concatenated after the quantum measurement. In addition, the concatenation operation used here is a simple vector operation that does not involve any learnable parameters.
For completeness, in a multi-head setting, where multiple target nodes $\bm{\nu}:=\{v_{k}\}_{k=1}^{K}$ are considered, the output embedding can either be fed into a QNN and processed as a quantum analogue of a dense layer or again it can be combined classically through a simple concatenation, yielding
$\mathbf{h}^{(L)}_{\bm{\nu}}(\bm{\theta})=\sigma\left(\left[\bigg\|_{k=1}^{K}\mathbf{h}_{v_{k}}^{(L)}(\bm{\theta}_{v})\right]\right),$ (15)
where $\bm{\theta}:=\{\bm{\theta}_{v_{k}}\}_{k=1}^{K}$. For simplicity, we assume target nodes have the same number of neighbors and that all QGCN circuits (or Ansätze) have the same depth $L$ across hops.
To this point, our analysis has focused on the unsupervised node-level task within the graph. By repeatedly performing message passing until the designated target node is reached, the resulting embedding captures both the intrinsic features of the target node and its topological relations within the graph. This final embedding yields the predicted graph-level output $\bar{y}(\bm{\theta}_{v})$, where $\bm{\theta}_{v}$ denotes the trainable parameters of the QGCN circuit. This formulation then enables supervised learning by comparing the model’s prediction with the true target value $y$. The loss function $\mathcal{L}$ quantifies the training error, and the parameters $\bm{\theta}_{v}$ are optimized by minimizing $\mathcal{L}(\bar{y}(\bm{\theta}_{v}),y)$, expressed as
$\bm{\theta}^{*}_{v}$ $=\arg\min_{\bm{\theta}_{v}}\mathcal{L}(y(\bm{\theta}_{v}),y).$ (16)
The parameters are optimized using classical gradient-based methods. Successful training yields a negligible prediction error, i.e. $|\bar{y}(\bm{\theta}_{v})-y|\approx 0$.
VI Results
We now present the training and testing results of our QGNN framework on the QM9 dataset, comparing its performance with that of classical GNNs across molecules of varying sizes. Here, we use GNN to refer specifically to the GraphSAGE model.
VI.1 Implementation details
Performance is evaluated using the Smooth L1 Loss and the R2 score, with an initial focus on a small subset of molecular samples. The selected Loss is chosen for its stability and robustness to outliers, making it particularly suitable for small sample sizes with high variability. Training employs the Adam optimizer [adam_2014] ($\beta_{1}=\beta_{2}=0.9$) with a decaying learning rate.
Molecules are processed in two stages: (i) an unsupervised atom-level aggregation, where atomic features are combined according to the molecular adjacency matrix to construct a global embedding of the molecule, and (ii) a supervised molecule-level stage, where this embedding is used to predict a target property that characterizes the entire molecule. Here, the loss function compares this prediction with true targets, guiding the training of the model to capture both local and global structural information. Algorithm 1 provides a high-level description of these steps. For the quantum models, the FM and the Ansatz are integrated directly into the model $\mathcal{M}$ definition.
Algorithm 1 (Q)GNNs on the QM9 dataset
1:Batch of molecules $G$ with adjacencies AA, atom features $F$, targets $T$, model $\mathcal{M}$
@@ VI.1 Implementation details
4:for each mol $(A,F,T)$ in batch do
5:  $prev\_out\leftarrow\mathbf{0}$, $atom\_outs\leftarrow[\,]$
6:  for each node $v$ in $G$ do
7:   $x\leftarrow\texttt{concat}(F[A[v]],prev\_out)$
8:   $out\leftarrow\mathcal{M}(x)$
9:   $prev\_out\leftarrow out$,
10:   Append outout to atom_outsatom\_outs
11:  end for
12:  $mol\_out\leftarrow\texttt{mean}(atom\_outs)$
13:  Append mol_outmol\_out to $T_{\mathrm{pred}}$
14:  Append $T$ to $T_{\mathrm{true}}$
15:  $loss\mathrel{+}=\texttt{criterion}(mol\_out,T)/|\texttt{batch}|$
16:end for
17:return $T_{\mathrm{true}}$, $T_{\mathrm{pred}}$, lossloss
In the proposed algorithm, all hops $k$ up to the number of atoms are considered; for instance, a 9-atom molecule undergoes 9 hops, with embeddings propagated from lower- to higher-index atoms, yielding a single-branch computational graph with one final head. Classical GNNs are implemented in PyTorch [pytorch], while QGNNs use Qadence [qadence2025]. The number of inputs/qubits are 8, matching the 7 node feature dimension plus one (accounting for the mean self-embedding of the target node) [^1] Features such as atomic number, chirality, degree, formal charge, radical electrons, hybridization, and scaled mass are kept, while aromaticity, hydrogen count, ring membership, and valence are excluded as they are consistently zero in the selected samples.. QM9 data are encoded via the Fourier Feature Map. Measurements assess local magnetization with tunable parameters $\bm{\omega}:=\{\omega_{i}\}_{i=1}^{N}$, and the average total magnetization is defined as an expectation value
$\langle{\mathcal{O}}\rangle=\frac{1}{N}\sum_{i=1}^{N}\omega_{i}\langle Z^{(i)}\rangle.$ (17)
For the first set of results, and since our primary goal is to demonstrate that a single quantum circuit, namely a QGCN circuit such as the one shown in Fig. 5, can handle graphs with varying numbers of nodes, we adopt a minimal quantum circuit design in which the QGCN serves as the sole aggregation mechanism. For each result, the same fixed-architecture circuit is applied at every aggregation step. Its structure (circuit depth) and number of input qubits (circuit width) remain constant within each implementation, even as the number of nodes varies across graphs, leading to a different number of nodes being input to the aggregator at each hop. Despite our chosen circuit design, the QGNN framework demonstrates reasonable training performance and, notably, stronger generalization on unseen data relative to its classical counterpart.
The hyperparameters of the quantum and classical models used in the results are detailed in Table 1. Unless otherwise specified, the models hyperparameters and evaluation metrics: Smooth L1 Loss and $R^{2}$ score, remain unaltered. Furthermore, the classical model is constructed to have a number of learnable parameters comparable to that of the quantum model, ensuring a fair comparison between the two frameworks.
samples FW inp/qub FM LR rr/layers params
Case 1 30 GNN 8 - 0.01 $[9,2]$ 284
30 QGNN 8 Fourier 0.01 $[3,5]$ 293
Case 2 30 GNN 8 - 0.03 $[8,4]$ 305
30 QGNN 8 Fourier 0.03 $[3,3,3]$ 293
Table 1: The table reports the hyperparameters of the classical GNN and quantum QGNN frameworks (FWs) for Case 1 and Case 2. For each configuration, we list the number of training samples, the number of classical input features (inp) or qubits (qub), the feature map type (FM, quantum models only), the initial learning rate (LR), the hidden-layer architecture for classical models or the circuit depth configuration $r$ for quantum models, and the total number of trainable parameters (params).
VI.2 Comparison Between GNNs and QGNNs